# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins – Dataset and Benchmarks

### v2

@@ Appendix F License and Data Usage
We accessed the Semantic Trails Dataset [29] via Figshare: https://doi.org/10.6084/m9.figshare.7429076.v2. The dataset is licensed under the Creative Commons CC0 1.0 license (https://creativecommons.org/publicdomain/zero/1.0/), which allows unrestricted copying, modification, and redistribution for any purpose, including commercial use, without requiring permission.
We accessed the Foursquare Open Source Places dataset via Hugging Face: https://huggingface.co/datasets/foursquare/fsq-os-places. Foursquare Open Source Places is licensed under the Apache License, Version 2.0:
Copyright 2024 Foursquare Labs, Inc. All rights reserved. Licensed under the Apache License, Version 2.0 (the "License"); you may not use this file except in compliance with the License. You may obtain a copy of the License at:  http://www.apache.org/licenses/LICENSE-2.0 Unless required by applicable law or agreed to in writing, software distributed under the License is distributed on an "AS IS" BASIS, WITHOUT WARRANTIES OR CONDITIONS OF ANY KIND, either express or implied. See the License for the specific language governing permissions and limitations under the License.
More details are available in Foursquare’s documentation: https://docs.foursquare.com/data-products/docs/access-fsq-os-places.
We released our Massive-STEPS dataset under the same Apache Version 2.0 License, and have included Foursquare Open Source Places’ license in our hosted dataset’s README file.

### v3

@@ Appendix G License and Data Usage
We accessed the Semantic Trails Dataset (Monti et al., 2018) via Figshare: https://doi.org/10.6084/m9.figshare.7429076.v2. The dataset is licensed under the Creative Commons CC0 1.0 license (https://creativecommons.org/publicdomain/zero/1.0/), which allows unrestricted copying, modification, and redistribution for any purpose, including commercial use, without requiring permission.
We accessed the Foursquare Open Source Places dataset via Hugging Face: https://huggingface.co/datasets/foursquare/fsq-os-places. Foursquare Open Source Places is licensed under the Apache License, Version 2.0. More details are available in Foursquare’s documentation: https://docs.foursquare.com/data-products/docs/access-fsq-os-places:
Copyright 2024 Foursquare Labs, Inc. All rights reserved. Licensed under the Apache License, Version 2.0 (the ”License”); you may not use this file except in compliance with the License. You may obtain a copy of the License at:  http://www.apache.org/licenses/LICENSE-2.0 Unless required by applicable law or agreed to in writing, software distributed under the License is distributed on an ”AS IS” BASIS, WITHOUT WARRANTIES OR CONDITIONS OF ANY KIND, either express or implied. See the License for the specific language governing permissions and limitations under the License.
We release our Massive-STEPS dataset under the same Apache Version 2.0 License, and have included Foursquare Open Source Places’ license in our hosted dataset’s README file.